OpenStudy (anonymous):

32% of what number is 51.2

OpenStudy (anonymous):

0.32x=51.2 x= 51.2/0.32 x=160

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

u r w c

OpenStudy (anonymous):

P=b*r b=P/r b=51.2/0.32 b=160,

OpenStudy (anonymous):

P is percentage, b is base and r is rate.

OpenStudy (anonymous):

thanks. just wanted to make sure the same formula applied. much appreciated!

OpenStudy (anonymous):

okay,,

OpenStudy (anonymous):

Trick?

OpenStudy (anonymous):

im looking to you, your here

OpenStudy (anonymous):

yes mam?